At the point of fission, a nucleus of 235U that has 92 protons is divided into two smaller spheres, each of which has 46 protons and a radius of 5.74 ✕ 10-15 m. What is the magnitude of the repulsive force pushing these two spheres apart? N
Answer #1

Electrostatic force is given by:

F = k*q1*q2/R^2

k = 9*10^9

q1 = q2 = +46e

e = 1.6*10^-19 C

R = distance between both sphere = 2*r

r = radius of each sphere = 5.74*10^-15 m

R = 2*5.74*10^-15 m

So Using these values:

F = 9*10^9*(46*1.6*10^-19)^2/(2*5.74*10^-15)^2

F = 3699.2 N

Please Upvote.